If NH3 (ammonia, pKb=4.74) is added to a solution of Fe(NO3)3 such that right after the addition we have 50 mL of 5.0 M NH3 and 0.0040 M Fe(NO3)3 (right after addition)
Find Kc for
Fe2+(aq) + 2OH-(aq) ---> Fe(OH)2 (s)

any type of explanation would be helpful please!
Answer #1

Here, [Fe3+] = [Fe2+] = 0.004 M or 4*10-3 M

pOH = 1/2 (pKb - Log[NH3]) = 1/2 (4.74 - Log5) = 2.0205

i.e. [OH-] = 10-2.0205 = 9.54*10-3 M

Now, Kc = 1/[Fe2+][OH-] = 1/(4*10-3 * 9.54*10-3) = 2.621*104